# Machine Unlearning: A Comprehensive Survey

### v1

@@ Abstract.
As the right to be forgotten has been legislated worldwide, many studies attempt to design unlearning mechanisms to protect users’ privacy when they want to leave machine learning service platforms. Specifically, machine unlearning is to make a trained model to remove the contribution of an erased subset of the training dataset. This survey aims to systematically classify a wide range of machine unlearning and discuss their differences, connections and open problems. We categorize current unlearning methods into four scenarios: centralized unlearning, distributed and irregular data unlearning, unlearning verification, and privacy and security issues in unlearning. Since centralized unlearning is the primary domain, we use two parts to introduce: firstly, we classify centralized unlearning into exact unlearning and approximate unlearning; secondly, we offer a detailed introduction to the techniques of these methods. Besides the centralized unlearning, we notice some studies about distributed and irregular data unlearning and introduce federated unlearning and graph unlearning as the two representative directions. After introducing unlearning methods, we review studies about unlearning verification. Moreover, we consider the privacy and security issues essential in machine unlearning and organize the latest related literature. Finally, we discuss the challenges of various unlearning scenarios and address the potential research directions.
machine unlearning
@@ 1. Introduction
Over the past decade, enormously increased data and fast hardware improvement have driven machine learning developments quickly. Machine learning (ML) algorithms and artificial intelligence (AI) are embedded into day-to-day applications and wearable devices (Mahdavinejad et al., 2018). It further collects more and more users’ information, including users’ private data, such as driving trajectory, medical and online shopping records. On the one hand, such an enormous amount of data helps to improve ML and AI development further. However, on the other hand, it threatens users’ privacy and demands the enormous requirement of data management to ensure information security and privacy in machine learning.
@@
Machine unlearning is recently proposed to protect users’ privacy when they want their data forgotten by a trained machine learning model. It draws increasing research attention because Right to be forgotten was legislated recently in many countries, such as the General Data Protection Regulation (GDPR) (Mantelero, 2013) in the European Union, PIPEDA privacy legislation in Canada (of the Privacy Commissioner of Canada, 2018), and the California Consumer Privacy Act (de la Torre, 2018) in the United States. The legislation requires that companies must take reasonable measures to achieve the erasure of personal data concerning. For a machine learning application, it means that individual users have the right to ask companies to delete their private data once collected from them. The deletion is not only erasing their data from a database. It needs to remove the contribution of the erased subset of whole training data from the already trained model, which process was first conceptualized as machine unlearning (Cao and Yang, 2015). Specifically, if a user (Alice) wants to exercise her right to be forgotten (Mantelero, 2013) (e.g., quitting a ML application), then the trained model of such application must ”unlearn” her data. Such a process includes two steps: first, a subset of the whole dataset previously used for training a ML model is requested to be erased; second, the ML model provider removes the contribution of the erased data from the trained model. A naive data-erasing method is retraining a new model from scratch (Bourtoule et al., 2021). However, the computation and storage costs of retraining are expensive, especially in complex learning tasks.
@@
Many researchers tried to find efficient and effective methods to realize unlearning rather than naive retraining, and there are several common challenges, which are summarized as follows. Stochasticity of training: A huge amount of randomness exists in the training process in machine learning, especially in complicated models’ training periods such as CNNs (Gu et al., 2018) and DNNs (Sze et al., 2017). This randomness makes the training result non-deterministic (Bourtoule et al., 2021) and raises difficulties for machine unlearning to estimate the contribution of the specific erased data sample. Incrementality of training: The training process in machine learning is an incremental procedure. In other words, the model update on the former data point will influence the contribution of the data points fed into the model after this data. Determining a way to effectively remove the influence of the to-be-erased data samples from the trained model is challenging for machine unlearning (Koh and Liang, 2017). Catastrophe of unlearning: Nguyen et al. (Nguyen et al., 2020) indicated that an unlearned model usually performs worse than the model retrained from scratch. Especially when an unlearning method tries to remove a huge amount of data samples, the degradation could be exponential. They referred to such sharp degradation as catastrophic unlearning (Nguyen et al., 2020). Although several studies have explored ways to mitigate unlearning degradation by bounding the loss function or restricting unlearning threshold, how to eliminate catastrophic unlearning is still an open problem. Recently, many studies put efforts into solving these three main challenges and proposed many novel mechanisms which promote the progress of machine unlearning.
@@
Machine UnlearningCentralized UnlearningUnlearning CategoriesExact Unlearning: (Cao and Yang, 2015; Bourtoule et al., 2021; Golatkar et al., 2021; Graves et al., 2021; Garg et al., 2020; Brophy and Lowd, 2021; Yan et al., 2022; Chen et al., 2022)Approximate Unlearning: (Wu et al., 2020, 2022a; Guo et al., 2020; Sekhari et al., 2021; Neel et al., 2021; Schelter et al., 2021; Ginart et al., 2019; Golatkar et al., 2020a; Mehta et al., 2022; Li et al., 2021; Warnecke et al., 2021; Nguyen et al., 2020, 2022a; Fu et al., 2022; Khan and Swaroop, 2021) Unlearning TechniquesSplit Unlearning: (Cao and Yang, 2015; Bourtoule et al., 2021; Golatkar et al., 2021; Graves et al., 2021; Garg et al., 2020; Brophy and Lowd, 2021; Yan et al., 2022; Chen et al., 2022)Certified Data Removal: (Wu et al., 2020, 2022a; Guo et al., 2020; Sekhari et al., 2021; Neel et al., 2021; Schelter et al., 2021; Ginart et al., 2019; Golatkar et al., 2020a; Mehta et al., 2022; Li et al., 2021; Warnecke et al., 2021)Bayesian-based Unlearning: (Nguyen et al., 2020, 2022a; Fu et al., 2022)Distributed and Irregular Data UnlearningFederated Unlearning: (Wang et al., 2022; Liu et al., 2021, 2022a; Gupta et al., 2021; Wu et al., 2022b)Graph Unlearning: (Chien et al., 2022; Chen et al., 2021a; Cong and Mahdavi, 2022)Unlearning Verification: (Hu et al., 2022; Thudi et al., 2022; Ma et al., 2022; Sommer et al., 2022; Huang et al., 2021a) Privacy and Security Issues in Machine UnlearningPrivacy Threats in UnlearningMembership Inference: (Chen et al., 2021b; Zanella-Béguelin et al., 2020; Shokri et al., 2017; Lu et al., 2022; Huang et al., 2021a; Graves et al., 2021; Golatkar et al., 2020b)Privacy Reconstruction: (Gao et al., 2022; Chen et al., 2021b; Zanella-Béguelin et al., 2020; Shintre et al., 2019; Baumhauer et al., 2022; Graves et al., 2021; Golatkar et al., 2020b)Unlearning ApplicationsAnomaly Removal: (Wang et al., 2019; Liu et al., 2022b; Cao et al., 2018; Du et al., 2019; Shibata et al., 2021; Ye et al., 2022; Ganhör et al., 2022)Data Unlearnable: (Huang et al., 2021b)
@@
Figure 1. Our taxonomy for machine unlearning. The introduction order will also follow this figure. We classify current unlearning literature into four main scenarios: centralized unlearning, distributed and irregular data unlearning, unlearning verification, and privacy and security issues in machine unlearning.
@@
The goal of this survey is to classify and systematize machine unlearning according to the reseaching problems and goals in unlearning processes, and to review their differences and connections, advantages and disadvantages. We first catalogue unlearning methods into four main parts, centralized unlearning, distributed and irregular data unlearning, unlearning verification, and privacy and security issues in machine unlearning, as shown in Fig. 1. Since there are many studies targeted at centralized unlearning, we use two sections to introduce related work in this scenario. We briefly classify the related unlearning methods into two main kinds, Exact unlearning (Cao and Yang, 2015; Bourtoule et al., 2021) and Approximate unlearning (Chien et al., 2022; Nguyen et al., 2020), where exact unlearning extends the idea of naive unlearning to design efficient retraining scheme with less computation cost. Approximate unlearning aims to unlearn a posterior that is approximate to the model retrained on the remaining dataset, only using the original learned model and the erased data. We then detailedly introduce the popular implementing techniques of these two unlearning kinds. In addition to these centralized unlearning methods, we introduce the distributed and irregular data unlearning scenarios, including federated unlearning (Wang et al., 2022; Liu et al., 2021), and graph unlearning (Chien et al., 2022; Chen et al., 2021a). After implementing unlearning, verifying the unlearning effect is equally important and draws lots of research attention. Therefore, we review the studies about unlearning verification subsequently. Moreover, we consider the privacy and security issues essential in machine unlearning, and we organize and review the related unlearning literature about privacy threats, defences, and unlearning applications in the survey.
@@
There is no comprehensive survey about machine unlearning because it is a pretty new research domain. Only several reviews on the specific unlearning scenarios were published on Arxiv. An introduction to machine unlearning, about the exact and approximate unlearning problems and how they are defined and solved by recently proposed methods, see (Mercuri et al., 2022). About certifiable machine unlearning for linear models, the algorithm introduction and experiments analysis can be found in (Mahadevan and Mathioudakis, 2021). Nguyen et al. (Nguyen et al., 2022b) summarized the general unlearning framework and added the unlearning verification part to it. They paid most of their attention to introducing problem formulations and technical definitions. None of them noticed and reviewed the privacy and security issues in machine unlearning. Moreover, they have not comprehensively discussed the differences and connections between different unlearning scenarios.
In our treatment, we review not only centralized machine unlearning but also other scenarios, including distributed and irregular data unlearning, unlearning verification, and privacy and security issues in machine unlearning. We discuss these unlearning scenarios in the literature and compare the difference and connections from each other. For all four unlearning scenarios, we list open questions that, if solved, would increase our understanding of machine unlearning or enhance our ability to design a suitable unlearning mechanism for different applications.
The main contributions of the paper are as follows.
@@
  • We systematically catalogue machine unlearning studies according to their scenarios, which can be briefly classified into, centralized unlearning, distributed and irregular data unlearning, unlearning verification, and privacy and security issues in machine unlearning.
@@
  • For the centralized unlearning scenario, which draws the most unlearning research attention, we further divide related studies into exact and approximate unlearning and illustrate the connections, pros, and cons among these works.
@@
  • We further explore the privacy threats that machine unlearning brings and discuss the applications of unlearning in defending against traditional security and privacy issues.
@@
  • We discuss different challenges between various machine unlearning scenarios, from centralized unlearning to federated unlearning, graph unlearning, and privacy and security issues in unlearning. We list the related open questions of each scenario and present potential research directions to solve them.
@@
The rest of the survey is organized as follows. The background of machine unlearning and the related technical tools are summarized in Section 2. After introducing the background, we will introduce the survey following the taxonomy order in Fig. 1. In Section 3, we present the primary categories of centralized unlearning: exact unlearning and approximate unlearning. In Section 4, techniques of centralized unlearning in different categories introduced in Section 3 are discussed in-depth. We collect and introduce the distributed and irregular data unlearning scenarios that we noticed, which are federated unlearning and graph unlearning, in Section 5. Section 6 introduces the unlearning evaluation and verification methods. The privacy and security issues in machine unlearning are divided into two sections. Section 7 introduces the privacy threats accompanied by machine unlearning. In Section 8, we discuss machine unlearning applications, which are mainly applied in dealing with security issues. In Section 9, we discuss the challenges of current unlearning methods in a general way. After that, we list the open gaps that new studies can target. At last, we summarise the survey in Section 10.
@@ 2.1. Machine Unlearning
There are many urgent demands that motivate the hot machine unlearning research. The first is the privacy-preserving demand. Since the right to be forgotten has been legislated worldwide, machine unlearning is often formulated as a privacy-preserving problem where users can ask for their data erasure from the ML models. Besides privacy reasons, many other reasons promote the development of machine unlearning. For example, one common reason is usability. During the long time usage period of a recommendation application, users’ information will sometimes change, such as moving changes in their location. The application will produce unsuitable recommendations if it cannot completely delete the incorrect data from the model. Another reason is security. Deep learning models are vulnerable to adversarial attacks (Yuan et al., 2019) such as data poisoning (Tian et al., 2022). Hence, how to detect and remove adversarial poisoned data is essential to ensure models’ security. Once the attack type is detected, the model must erase these adversarial data’s influence using the unlearning mechanism.
Figure 2. Machine unlearning process.
@@
To better understand how unlearning mechanisms work, we first introduce the unlearning problem and process following the machine unlearning framework demonstrated in Fig. 2. Let $\mathcal{Z}$ denote as a space of data items, then the particular (full) training dataset is $D\in\mathcal{Z}$. A learning process can be demonstrated as Step 0 in Fig. 2, i.e., training a model $M$ using an algorithm $\mathcal{A}$ on the training dataset $D$, denoted as $M=\mathcal{A}(D)$, where model $M$ is in a hypothesis space $\mathcal{H}$.
@@
The unlearning process begins at an unlearning request when a user wants to remove his data $D_{e}$ from the trained model, Step 1 in Fig. 2. The requested unlearning data $D_{e}$ can be data samples, classes or graph nodes. Then, in Step 2, the server removes the contribution of $D_{e}$ using a machine unlearning algorithm $\mathcal{U}$. The unlearned model can be expressed as $M_{D\backslash D_{e}}=\mathcal{U}(D,D_{e},\mathcal{A}(D))$. In dealing with an unlearning request, we hope that the unlearned model $\mathcal{U}(D,D_{e},\mathcal{A}(D))$ is the same as the retrained model $\mathcal{A}(D\backslash D_{e})$ (i.e., $\mathcal{U}(D,D_{e},\mathcal{A}(D))\simeq\mathcal{A}(D\backslash D_{e})$). Most unlearning studies ended at this step. However, we do need an effective evaluation metric to assess the unlearning effect as Step 3 in Fig.2. Therefore, we add the verification as the last step in the unlearning process, and we will introduce the related literature later.
2.2. Unlearning Evaluation Metrics

### v2

@@ Abstract.
As the right to be forgotten has been legislated worldwide, numerous studies have sought to design unlearning methods to protect users’ privacy when they want to remove their data from machine learning service platforms. This has given rise to the concept of “machine unlearning” — a field dedicated to removing the influence of specified samples from trained models. This survey aims to systematically classify a wide range of machine unlearning studies, discussing their differences, connections, and open problems. We categorize current unlearning methods into four key areas: centralized unlearning, federated unlearning, unlearning verification, and privacy and security issues in unlearning. Given that centralized unlearning is the primary focus, we further divide its discussion into two parts: first, at a higher level, we classify centralized unlearning into exact unlearning and approximate unlearning; second, we provide a detailed introduction to the techniques used in these methods. Next, we introduce federated unlearning, another emerging area that has garnered significant attention. Following the discussion on unlearning methods, we review studies on unlearning verification and audit, which assess the effectiveness of existing unlearning algorithms. Additionally, we consider the privacy and security issues essential in machine unlearning and compile the related literature. Finally, we discuss the challenges of various unlearning scenarios and highlight potential research directions.
Federated Unlearning; Machine Unlearning; Unlearning Verification
@@ 1. Introduction
Over the past decade, enormously increased data and fast hardware improvement have driven machine learning developments quickly. Machine learning (ML) algorithms and artificial intelligence (AI) are embedded into day-to-day applications and wearable devices (Mahdavinejad et al., 2018). It continuously collects increasing amounts of user information, including private data such as driving trajectories, medical records, and online shopping histories (Kirkpatrick et al., 2017; Rolnick et al., 2019). On the one hand, such an enormous amount of data helps to further advance ML and AI development. On the other hand, however, it poses a threat to users’ privacy and creates a significant need for robust data management to ensure information security and privacy in ML (Tian et al., 2022).
@@
Machine unlearning has drawn growing research attention as the recent legislation of the ”Right to be Forgotten” in many countries. Notable instances include the GDPR in the European Union (Mantelero, 2013), the PIPEDA privacy legislation in Canada (of the Privacy Commissioner of Canada, ), and the California Consumer Privacy Act in the United States (de la Torre, 2018). According to these laws, companies must take reasonable measures to guarantee that personal data is deleted upon request. It indicates that individual users have the right to request companies to remove their private data, which was previously collected for ML model training. The deletion is not only erasing their data from a database, but it also needs to delete the influence of the specified samples from trained models. The process of data removal from models was first conceptualized as machine unlearning (Cao and Yang, 2015). Specifically, suppose a user (Alice) wants to exercise her right (Mantelero, 2013) when quitting a ML application, then the trained model of such application must ”unlearn” her data. Such a process includes two steps: first, a subset of the dataset previously used for ML model training is requested to be deleted; second, the ML model provider erases the contribution of these data from the trained models. A naive data-erasing method is retraining a new ML model from scratch (Bourtoule et al., 2021). However, the computation and storage costs of retraining are expensive, especially in complex learning tasks.
@@
Many researchers have tried to find efficient and effective methods to implement unlearning rather than naive retraining, and there are several common challenges, which are summarized as follows. (1) Stochasticity of training: A huge amount of randomness exists in the training process in machine learning, especially in complicated models’ training periods such as CNNs (Gu et al., 2018) and DNNs (Sze et al., 2017). This randomness makes the training results non-deterministic (Bourtoule et al., 2021) and raises challenges for machine unlearning to estimate the impact of the typical erased samples. (2) Incrementality of training: The training process in machine learning is incremental, meaning that the model update from one data point influences the contribution of subsequent data points fed into the model. Deciding a way to effectively remove the contributions of the to-be-erased samples from the trained model is challenging for machine unlearning (Koh and Liang, 2017). (3) Catastrophe of unlearning: Nguyen et al. (Nguyen et al., 2020) indicated that an unlearned model typically has worse model utility than the model retrained from scratch. The degradation would be exponential, especially when a method tries to delete a huge amount of data samples. They referred to such sharp degradation as catastrophic unlearning (Nguyen et al., 2020). Although several studies mitigate model utility degradation by bounding the loss function or restricting the unlearning update threshold, eliminating catastrophic unlearning remains an open problem. Recently, many studies have put efforts into solving these three main challenges and proposed many novel mechanisms that promote the progress of machine unlearning.
@@
Machine UnlearningCentralized UnlearningUnlearning CategoriesExact Unlearning: (Cao and Yang, 2015; Bourtoule et al., 2021; Golatkar et al., 2021; Graves et al., 2021; Garg et al., 2020; Brophy and Lowd, 2021; Yan et al., 2022; Chen et al., 2022a; Schelter et al., 2023; Koch and Soll, 2023; Hu et al., 2024a)Approximate Unlearning: (Wu et al., 2020, 2022a; Guo et al., 2020; Sekhari et al., 2021; Neel et al., 2021; Schelter et al., 2021; Ginart et al., 2019; Golatkar et al., 2020a; Mehta et al., 2022; Li et al., 2021; Warnecke et al., 2023; Nguyen et al., 2020, 2022a; Fu et al., 2022; Khan and Swaroop, 2021; Tarun et al., 2023; Wang et al., 2023a; Chundawat et al., 2023; Warnecke et al., 2024; Liu et al., 2023a; Lin et al., 2023; Cha et al., 2024) Unlearning TechniquesSplit Unlearning: (Cao and Yang, 2015; Bourtoule et al., 2021; Golatkar et al., 2021; Graves et al., 2021; Garg et al., 2020; Brophy and Lowd, 2021; Yan et al., 2022; Chen et al., 2022a)Certified Data Removal: (Wu et al., 2020, 2022a; Guo et al., 2020; Sekhari et al., 2021; Neel et al., 2021; Schelter et al., 2021; Ginart et al., 2019; Golatkar et al., 2020a; Mehta et al., 2022; Li et al., 2021; Warnecke et al., 2023)Bayesian-based Unlearning: (Nguyen et al., 2020, 2022a; Fu et al., 2022)Graph Unlearning: (Chien et al., 2022a; Chen et al., 2022b; Li et al., 2024; Cong and Mahdavi, 2022; Wang et al., 2023b; Wu et al., 2023a; Zhang et al., 2024; Wu et al., 2023b)Federated (Distributed) Unlearning: (Wang et al., 2022a, 2024; Fraboni et al., 2024; Wang et al., 2023c; Liu et al., 2021, 2022a; Zhang et al., 2023a; Gupta et al., 2021; Wu et al., 2022b)Unlearning Verification: (Hu et al., 2022; Guo et al., 2023; Thudi et al., 2022a; Ma et al., 2022; Sommer et al., 2022; Huang et al., 2021a; Thudi et al., 2022b) Privacy and Security Issues in Machine UnlearningPrivacy Threats in UnlearningMembership Inference: (Chen et al., 2021; Zanella-Béguelin et al., 2020; Shokri et al., 2017; Lu et al., 2022; Huang et al., 2021a; Graves et al., 2021; Golatkar et al., 2020b; Hu et al., 2024b)Privacy Reconstruction: (Gao et al., 2022; Chen et al., 2021; Zanella-Béguelin et al., 2020; Shintre et al., 2019; Baumhauer et al., 2022; Graves et al., 2021; Golatkar et al., 2020b; Hu et al., 2024b; Zhang et al., 2023b)Security Threats in Unlearning : (Marchant et al., 2022; Zhao et al., 2024; Di et al., 2022; Hu et al., ; Liu et al., 2024)Unlearning ApplicationsAnomaly Removal: (Wang et al., 2019; Wei et al., 2023; Liu et al., 2022b; Cao et al., 2018; Du et al., 2019; Shibata et al., 2021; Ye et al., 2022; Ganhör et al., 2022)Data Unlearnable: (Huang et al., 2021b)
@@
Figure 1. Our taxonomy for machine unlearning. The introduction order will also follow this figure. We classify the current unlearning literature into four main scenarios: centralized unlearning, federated unlearning, unlearning verification, and privacy and security issues in machine unlearning.
@@
This survey aims to classify and systematize machine unlearning methods based on the research problems and objectives in the unlearning process and to review their differences and connections, as well as their advantages and disadvantages. We first categorize unlearning approaches into four main areas: centralized unlearning, federated unlearning, unlearning verification, and privacy and security issues in machine unlearning, as shown in Fig. 1. Since many studies focus on centralized unlearning, we dedicate two sections to introducing corresponding studies in this area. We classify these unlearning methods into: Exact unlearning (Cao and Yang, 2015; Bourtoule et al., 2021) and Approximate unlearning (Chien et al., 2022a; Nguyen et al., 2020), where exact unlearning extends the idea of naive unlearning to design efficient retraining scheme with less computation cost. Approximate unlearning tries to unlearn a model approximate to the model retrained using the remaining dataset. We also provide a detailed introduction to the popular techniques used to implement these two types of unlearning. In addition to centralized unlearning methods, we introduce federated (distributed) unlearning studies (Wang et al., 2022a; Liu et al., 2021), which are gaining increasing importance. After discussing unlearning implementation methods, verifying the unlearning effect is equally crucial and has garnered significant research attention (Thudi et al., 2022b). Therefore, we subsequently review related research on unlearning verification. Finally, we consider privacy and security issues in machine unlearning essential, organizing and reviewing the related publications on privacy and security threats, defenses, and unlearning applications.
@@
There are only a few surveys on machine unlearning because it is a relatively new research domain, and most of them are preprints on Arxiv, focusing on certain unlearning scenarios. For an introduction to machine unlearning, including discussions on exact and approximate unlearning problems and their solutions through recently proposed methods, see (Mercuri et al., 2022). For information on provable machine unlearning for linear models, including algorithm introductions and experimental analysis, refer to (Mahadevan and Mathioudakis, 2021). For an overview of federated unlearning, see (Liu et al., 2023b; Wang et al., 2023d), and for graph unlearning, see (Said et al., 2023). While Nguyen et al. (Nguyen et al., 2022b) summarized the general unlearning framework and added the unlearning verification part to it, they focused primarily on introducing problem formulations and technical definitions. However, none of these surveys reviewed the privacy and security issues in machine unlearning, nor did they comprehensively discuss the differences and connections between various unlearning scenarios.
Compared with existing surveys on machine unlearning, we performed a systematic literature review (SLR) following the guidelines presented by (Keele et al., 2007) and as exemplified in (Kubrak et al., 2022). The contributions of this paper include:
@@
  • We systematically catalog machine unlearning studies according to their scenarios, which can be briefly classified into centralized unlearning, distributed unlearning, unlearning verification, and privacy and security issues in machine unlearning.
@@
  • For the centralized unlearning scenario, which draws the most attention, we divide related studies into exact and approximate unlearning and illustrate the connections, pros, and cons among these works.
@@
  • We further explore the privacy and security threats that target machine unlearning and discuss the applications of unlearning in defending against traditional security and privacy issues.
@@
  • We discuss different challenges between various machine unlearning areas, from centralized unlearning to federated unlearning, unlearning verification, and privacy and security issues in unlearning. We list the related open questions of each scenario and present potential research directions to solve them.
@@
The survey’s remaining sections are arranged as follows. We show a systematic literature review model of the survey in Section 2. The basic knowledge of machine unlearning and related technical tools are summarized in Section 3. After introducing the background, we introduce the survey following the taxonomy order in Fig. 1. The primary content of centralized unlearning is presented in Section 4, which includes an introduction from two main unlearning categories and corresponding in-depth techniques. We collect and introduce the federated unlearning scenarios in Section 5. Section 6 introduces the unlearning evaluation and verification methods. The privacy and security issues in machine unlearning are divided into two sections. Section 7 introduces the privacy and security threats accompanied by machine unlearning. Section 8 discusses machine unlearning applications, which are mainly applied in dealing with security issues. In Section 9, we discuss the challenges of current unlearning methods in a general way. After that, we enumerate the unaddressed issues for further research. At last, in Section 10, we provide a summary of the survey.
2. The SLR Method of the Survey
We performed a systematic literature review (SLR) of the extant research on machine unlearning. The main goal was to compile a large body of research on machine unlearning and categorize them from various perspectives to facilitate analysis. Therefore, we employ the SLR method because it is especially appropriate for finding pertinent literature on a certain research topic (Keele et al., 2007). The main process of the SLR method for the survey is presented as follows.
We first designed a search string according to the review protocol (Keele et al., 2007), identified appropriate digital databases, and defined the data extraction strategy. We focused on the keywords “unlearning” and “machine unlearning” in the search string, formulated as “unlearning OR machine unlearning”. We used this search string in IEEE Xplore, ACM Digital Library, Scopus, and the Web of Science to find relevant papers. Additionally, we conducted a search on Arxiv to identify further relevant literature. Consequently, a total of 972 papers were retrieved (the search was conducted on 7 July 2024). We then limited the publication years to those after 2020 and ensured that the keyword “unlearning” appeared in the title and abstract. Moreover, after filtering out duplicates and papers with fewer than six pages, 261 papers remained. Referring to the Google Scholar top publications and the China Computer Federation (CCF) recommendations lists, we focused on reviewing 103 papers from top venues. After including 33 references for related techniques, we reviewed a total of 136 references.
During the systematic literature review process, the data extraction strategy is defined as follows. Besides the basic information of papers (title, author, publication venue, year), we first categorize papers into the four main classes as Fig. 1 according to the research problems. Then, we extract the techniques utilized, the evaluating metrics, and the datasets employed in these papers, which will all be introduced in the following context.
@@ 3.1. Machine Unlearning
There are several urgent demands driving machine unlearning research. The foremost is the demand for privacy preservation. With the “right to be forgotten” being legislated globally (Mantelero, 2013; of the Privacy Commissioner of Canada, ), machine unlearning ensures that users can request the erasure of their data from trained ML models. In addition to privacy concerns, other factors are promoting the development of machine unlearning. One significant factor is model utility. In the real world, vast amounts of data are generated daily, necessitating prompt updates to model services, as outdated data can negatively impact model performance (Wang et al., 2022b; Rolnick et al., 2019; Lopez-Paz and Ranzato, 2017). An effective unlearning mechanism is crucial for mitigating the adverse effects of outdated and incorrect data on model utility. Another critical factor is security. Adversarial attacks and data poisoning (Tian et al., 2022) can easily compromise deep learning models. Therefore, detecting and removing adversarial and poisoned data is essential to ensure model security. Once an attack type is identified, the model must erase the influence of these adversarial data using the unlearning mechanism.
@@
To better understand how unlearning mechanisms work, we first introduce the unlearning problem and process following the machine unlearning framework demonstrated in Fig. 2, and we summarize the notations in Table 1. Let $\mathcal{Z}$ denote as a space of data items; the particular (full) training dataset is $D\in\mathcal{Z}$. A learning process can be demonstrated as Step 0 in Fig. 2, i.e., training a model $M$ using an algorithm $\mathcal{A}$ on the training dataset $D$, denoted as $M=\mathcal{A}(D)$, where model $M$ is in a hypothesis space $\mathcal{H}$.
@@
The unlearning process begins at an unlearning request when a user wants to erase his specified data $D_{e}$ from the trained model, Step 1 in Fig. 2. The requested unlearning data $D_{e}$ can be data samples (Bourtoule et al., 2021), classes (Warnecke et al., 2023), or graph nodes (Chen et al., 2022b). Then, in Step 2, the server removes the contribution of $D_{e}$ using a machine unlearning algorithm $\mathcal{U}$. The unlearned model can be described as $M_{D\setminus D_{e}}=\mathcal{U}(M,D,D_{e})$. When dealing with an unlearning request, we hope that the unlearned model $\mathcal{U}(M,D,D_{e})$ is the same as the retrained model $\mathcal{A}(D\setminus D_{e})$ (i.e., $\mathcal{U}(M,D,D_{e})\simeq\mathcal{A}(D\setminus D_{e})$). Most unlearning studies ended at this step. However, we do need an effective evaluation metric to assess the unlearning effect, as in Step 3 in Fig. 2. Therefore, we add the verification step as the last step in the unlearning process, and we will introduce the related literature later.
Figure 2. Machine Unlearning Process
Table 1. Notations in Machine Unlearning
Symbols Description Symbols Description
$\mathcal{Z}$ data items space $\mathcal{H}$ model parameters space
$D=(X,Y)$ the training dataset with inputs $X$ and labels $Y$ $\mathcal{A}(\cdot)$ the ML algorithm
$D_{e}=(X_{e},Y_{e})$ the unlearned (erased) dataset $\mathcal{U}(\cdot)$ the unlearning algorithm
$D_{r}=D\setminus D_{e}$ the remaining dataset $M$ the model with parameters $\theta$
$D_{\text{probe}}$ the probing data used in attacking $M_{D\backslash D_{e}}$ the unlearned model
3.2. Machine Unlearning Evaluation Metrics